OpenStudy (anonymous):

Solve the equation: 8n - 2 + 4n = 66 - 5n

OpenStudy (anonymous):

n=11

OpenStudy (anonymous):

8n - 2 + 4n = 66 - 5n 12n - 2 = 66 - 5n 17n - 2 = 66 17n = 68 n = 4

OpenStudy (anonymous):

gabriellerencee is correct.

OpenStudy (anonymous):

Rosa34 I panicked for a second and had to recheck my answer!

OpenStudy (anonymous):

yes i did an did turns out i added wrong.